# The Elephant in the Room: Rethinking the Usage of Pre-trained Language Model in Sequential Recommendation

### v2

@@ Abstract.
Sequential recommendation (SR) has seen significant advancements with the help of Pre-trained Language Models (PLMs). Some PLM-based SR models directly use PLM to encode user historical behavior’s text sequences to learn user representations, while there is seldom an in-depth exploration of the capability and suitability of PLM in behavior sequence modeling. In this work, we first conduct extensive model analyses between PLMs and PLM-based SR models, discovering great underutilization and parameter redundancy of PLMs in behavior sequence modeling. Inspired by this, we explore different lightweight usages of PLMs in SR, aiming to maximally stimulate the ability of PLMs for SR while satisfying the efficiency and usability demands of practical systems. We discover that adopting behavior-tuned PLMs for item initializations of conventional ID-based SR models is the most economical framework of PLM-based SR, which would not bring in any additional inference cost but could achieve a dramatic performance boost compared with the original version. Extensive experiments on five datasets show that our simple and universal framework leads to significant improvement compared to classical SR and SOTA PLM-based SR models without additional inference costs.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
A recommender system is a type of information processing system that is deployed to forecast user preferences and provide suitable items or information. Its application spans various domains such as e-commerce, advertising, social media, and streaming services. Intensive studies about recommender systems of various real-world scenarios have found that historical interactions are important signals to model user preferences. Hence, sequential recommendation (SR) is a promising topic in the recommender systems community, aiming to capture the users’ dynamic preferences and predict the next item they like based on their historical behaviors. Early studies mainly adopt Markov chains to model the transition patterns of users (FPMC). In recent years, due to the rapid development of deep learning, different kinds of neural networks have been introduced into the SR, which resulted in notable enhancements. The most representative works include GRU4Rec (GRURec), Caser (Tang_Wang_2018), SR-GNN (SR-GNN), SASRec (SASRec), and BERT4Rec (BERT4Rec).
@@
With the thriving of PLMs, many researchers have begun exploring their potential application in SR. Various studies have utilized PLMs as item text encoders, aiming to enhance item representations with the rich prior knowledge embedded within PLMs (zhou2020s3; IDA-SR; TransRec; UniRec; yuan2023go). These methods always leverage a PLM to encode item text (e.g., title, brand) into text embeddings, which are then utilized to substitute or augment the original ID embeddings. Recently, propelled by the significant reasoning and long text modeling abilities demonstrated by Large Language Models (LLMs) in a variety of natural language processing tasks, some pioneering researchers have sought to harness the advanced sequence modeling capabilities of PLMs to enhance the performance of SR. One line of work directly uses fixed LLMs for recommendations via prompt or in-context learning, which achieve promising performance under the few-shot or zero-shot settings (gao2023chat; liu2023chatgpt). However, due to the significant disparity between language and behavior modeling, these methods often perform poorly compared to conventional ID-based SR models when the user’s historical interactions are sufficient. Inspired by this, recent works mostly adopt recommendation objectives to fine-tune PLMs for adaptation, which achieve promising results (geng2022recommendation; cui2022m6; zhang2023recommendation; bao2023tallrec; li2023text; qu2023thoroughly).
@@
Despite the demonstrated efficacy of PLM-enhanced methods in the existing literature, a critical question remains largely unexplored: do PLMs effectively and efficiently boost sequential recommendation by providing accurate prior knowledge and superior sequence modeling capabilities as theoretically expected? While some research has investigated the effect of PLMs in item representation learning (TransRec; yuan2023go; wei2023llmrec), their impact on behavior sequence modeling has yet to be thoroughly examined. Consequently, our study aims to figure out two questions: (1) whether the powerful sequence modeling capabilities of PLMs have been fully and economically utilized in the behavior sequence modeling of SR. (2) If not, is there a more effective, universal, and economical way for us to employ PLMs for SR?
@@
For question one, we conduct two motivating experiments (in Sec. 3) using a representative SOTA PLM-enhanced multi-domain SR model, RECFORMER (li2023text). The findings highlight two critical aspects: (1) RECFORMER functions differently from its base PLM backbone in behavior modeling. Its attention maps exhibit a clear functional stratification, with higher layers (more crucial for SR adaptation) more assimilated to conventional ID-based SR models. (2) There appears to be significant parameter redundancy when employing PLMs for behavior modeling. The global attentions of the sequence display considerable resemblance among different heads and layers in the same functional stratification. Only tuning 1/4 selected layers of RECFORMER can yield performance comparable to or even surpassing the original RECFORMER.
@@
Based on the above observation, we implement several enhanced variants of RECFORMER with simplified sequence modeling methods borrowed from classical ID-based SR models such as SASRec and BERT4Rec. We discover the key factor of improvement borrowed from PLM to SR: item representations built by behavior-tuned PLM. Experimental results indicate that PLM-involved SR models with simpler sequence models can achieve comparable results to RECFORMER with much better efficiency, suggesting that PLMs are not that suitable for behavior sequence modeling, for their superior sequence modeling and reasoning capacities in NLP are not fully leveraged and transferred into the task of SR. Consequently, we conclude a simple, effective, and universal framework to efficiently take advantage of PLM for SR: (a) adopting simplified (ID-based) sequential models for behavior sequence modeling (for question one), and (b) using behavior-tuned PLMs (rather than vanilla PLMs) for item embedding initialization (for question two). Further experiments demonstrate that our behavior-tuned PLMs initialized item embeddings are universal with different (simple or complicated) sequential models and settings. The contributions are concluded as follows:
@@
  • We introduce a novel framework that uses behavior-tuned PLMs for item initialization and simplified methods for behavior sequence modeling. Our framework is simple, effective, and universal, leading to significant improvements over PLM-based and classical SR baselines on different settings without additional inference costs.
@@ 2.1. Sequential Recommendation
Sequential recommendation (SR) is a widely researched task in the field of recommender systems, aiming to model the users’ dynamic behavior patterns based on their historical interactions. The very early works on sequential recommendation usually adopt Markov chains (MCs) to model the behavioral patterns of users. For instance, FPMC (FPMC) integrates MF and MCs to simultaneously model users’ overarching preferences and their historical interactions. Beyond first-order MCs, several approaches have embraced higher-order MCs, which incorporate a broader array of preceding items for sequential modeling (He_Kang_McAuley_2017; Ruining_Julian_2016). Later, with the development of deep learning methods, various deep neural networks have been introduced into SR to better capture users’ dynamic preferences. RNNs such as Gated Recurrent Unit (GRU) (GRU) are introduced to model sequential behaviors, including session-based GRU (e.g., GRU4Rec (GRURec)) and user-based GRU (Donkers_Loepp_Ziegler_2017)). Moreover, CNNs have also been demonstrated to effectively capture short-term sequential dynamics using both horizontal and vertical convolutional filters (Tang_Wang_2018). Many Graph Neural Network (GNN) based models (SR-GNN; gnn2022evolutionary; gnn2022exploiting; gnnsr), were also implemented by researchers to seek better performance. Recently, due to the notable achievements in sequence modeling made by Transformers (transformers) in the field of natural language processing, many attention-based models have also been intensively explored and adopted in sequential recommendation. SASRec (SASRec) employs a unidirectional Transformer to learn dynamic user preferences in an autoregressive pattern. BERT4Rec (BERT4Rec) utilizes bidirectional Transformers with masked item prediction as an additional training signal to better understand users’ historical interactions for precise recommendations. LSSA (LSSA) introduces a network architecture that combines long-term and short-term self-attention mechanisms, designed to comprehensively model users’ enduring preferences alongside their immediate needs. Self-supervised learning (SSL) is also introduced by recent works for improving the performance and training efficiency of SR models (qiu2022contrastive; chen2022intent; xie2022contrastive; wu2022selective; zhou2020s3).
Besides, some recent efforts explore the cross-domain sequential recommendation (ma2019pi; hao2021adversarial; cao2022contrastive), while most of them face challenges to smoothly transfer the generalized information to new domains. To address this, in addition to adopting Transformers for behavior modeling, some works (TransRec; UniRec; yuan2023go) further use PLMs and vision encoders to fuse modality information into item representations for a more comprehensive understanding of item contents, functioning as the bridge across different datasets.
@@ 2.2. PLM for Recommendation
With the thriving of foundation models, many applications of PLMs for SR have been researched, hoping to leverage the rich knowledge and powerful sequence modeling capabilities of PLMs to enhance SR. Some works propose to use them as text encoders (e.g., build item representations from their textual representations based on PLMs rather than merely using the conventional random initialized ID embeddings). For example, IDA-SR (IDA-SR) and UniSRec (UniRec) apply BERT to obtain item representations from their corresponding texts. LLM2BERT4Rec (LLM-SR) utilizes an embedding model with the titles of the items to retrieve their embeddings. Besides, some researchers try to convert users’ historical interactions and items into plain text and utilize them to fine-tune the PLMs for SR (qu2023thoroughly; M6-rec; li2023text; geng2022recommendation).
Recently, the emergence of Large Language Models (LLMs) (GPT3; gpt4; instruction-tuning) has revolutionized the field of natural language processing, greatly improving the performances on various natural language processing tasks with their powerful reasoning abilities and rich world knowledge. Therefore, some pioneer works explore the possibility of directly using powerful LLMs for recommendation via prompt or in-context learning (hou2023large; gao2023chat; wang2023zero; liu2023chatgpt; kang2023llms). Although these methods achieve impressive results in explanation-generation tasks and cold-start scenarios, they still perform poorly in SR when interactions are sufficient since behavioral information is essential for making precise recommendations. Another line of work considers introducing collaborative signals into LLMs to improve their recommendation performance. TallRec (bao2023tallrec) and InstructRec (zhang2023recommendation) construct instruction-tuning datasets and perform instruct-tuning on LLMs to adapt them into SR tasks. RLMRec (ren2023representation) aligns the semantic space of LLMs with collaborative relational signals through cross-view alignment to improve the quality of representations. BIGRec (bao2023bi) treats SR as a text generation task, generating item titles by a two-step approach with the LLM fine-tuned on user behavioral sequences. Nevertheless, these methods lack a thorough examination of whether PLMs’ capabilities are being effectively and efficiently harnessed. In our paper, we deeply explore the involvement of PLMs in SR tasks. We intend to scrutinize whether their formidable abilities are truly being utilized to their full potential and offer insights on how to make the most out of PLMs in the context of SR, both efficiently and effectively.
@@ 3. Analyses on Behavior Sequence Modeling of PLM-based SR Models
In this section, we comprehensively analyze the capability and suitability of PLM-based SR models on behavior sequence modeling. Without loss of generality, we employ the SOTA PLM-based SR model RECFORMER (li2023text) for the following explorations.
@@ 3.1. Background of RECFORMER
RECFORMER is a PLM-based SR model that comprises a classical PLM Longformer (longformer) and two additional embeddings that indicate the type of token (determining the token belongs to the special token [CLS], textual attribute key or textual attribute value) and its associated item (which item the token belongs to). It transforms items and behavioral sequences into text sequences and then utilizes the behavior-tuned model to obtain their corresponding representations for providing recommendations. RECFORMER’s training process has three steps, including pre-training and two-stage fine-tuning. To simplify, we designate these stages as stage-PT, stage-FT1, and stage-FT2 respectively. (1) In stage-PT, RECFORMER adopts Masked Language Modeling (MLM) and item-item contrastive (IIC) tasks to pre-train the model. Considering the computational costs, RECFORMER treats in-batch next items as negative instances instead of negative sampling or fully softmax for the IIC task. (2) In stage-FT1, RECFORMER encodes all items before each epoch starts, obtaining their corresponding item embeddings, then fine-tuning all parameters with these frozen embeddings (for efficiency considering the large size of item candidates). (3) In stage-FT2, the RECFORMER selects the optimal model from stage-FT1 to acquire its item embeddings and fix them. Stage-FT2 only tunes the parameters of the behavior sequence modeling part (similar to stage-FT1, but with the item embeddings remaining fixed and not updated each epoch). This strategy of maintaining fixed item embeddings is primarily motivated by the significant computational demands associated with generating all item representations during the cross-entropy loss calculation. Therefore, the fine-tuning process of RECFORMER could be simplified as: (1) obtaining the item embeddings from behavior-tuned Longformer, and (2) tuning PLM solely for behavior sequence modeling with fixed item embeddings. The latter step is our focus.
@@ 3.2. Analysis on Attention Discrepancy
Intuitively, the sequential patterns of items in SR and tokens in LM significantly differ. SR tends to prioritize the most recent interactions of users, focusing more on their immediate past behavior to make predictions. On the other hand, LM concentrates more on maintaining linguistic consistency and adhering to the syntactic structure of the text, ensuring coherence and grammatical accuracy throughout the content. To verify this assumption, we visualize the attention distributions of the special token [CLS] that is used to get the user representation for recommendation in RECFORMER (and its backbone Longformer (longformer) for comparisons).
Specifically, we follow the setting in (li2023text) to pre-train and fine-tune RECFORMER, and stack the attention scores from different heads in the same layer on the vertical axis. As for Longformer, we directly utilize its original checkpoint, drawing the attention maps in the same way as RECFORMER. It’s important to highlight that RECFORMER reverses the user’s interaction sequence. Consequently, within the attention maps, items positioned to the left symbolize those with which the user has most recently interacted. In Fig. 1, we provide the attention maps of two instances from Pantry and Instruments (amazon) with its length of historical sequence as three (for better visualization). Upon examining various instances from multiple domains, we have some universal findings:

### v3

@@ Abstract.
Sequential recommendation (SR) has seen significant advancements with the help of Pre-trained Language Models (PLMs). Some PLM-based SR models directly use PLM to encode user historical behavior’s text sequences to learn user representations, while there is seldom an in-depth exploration of the capability and suitability of PLM in behavior sequence modeling. In this work, we first conduct extensive model analyses between PLMs and PLM-based SR models, discovering great underutilization and parameter redundancy of PLMs in behavior sequence modeling. Inspired by this, we explore different lightweight usages of PLMs in SR, aiming to maximally stimulate the ability of PLMs for SR while satisfying the efficiency and usability demands of practical systems. We discover that adopting behavior-tuned PLMs for item initializations of conventional ID-based SR models is the most economical framework of PLM-based SR, which would not bring in any additional inference cost but could achieve a dramatic performance boost compared with the original version. Extensive experiments on five datasets show that our simple and universal framework leads to significant improvement compared to classical SR and SOTA PLM-based SR models without additional inference costs. Our code can be found in https://github.com/777pomingzi/Rethinking-PLM-in-RS.
@@
journalyear: 2024copyright: acmlicensedconference: 18th ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italybooktitle: 18th ACM Conference on Recommender Systems (RecSys ’24), October 14–18, 2024, Bari, Italydoi: 10.1145/3640457.3688107isbn: 979-8-4007-0505-2/24/10ccs: Information systems Recommender systems
@@ 1. Introduction
Recommender system is an information processing system that is deployed to forecast user preferences and provide suitable items or information. Its application spans various domains such as e-commerce, advertising, and streaming services. Intensive studies in various real-world scenarios have found that historical interactions are important signals to model user preferences. Hence, sequential recommendation (SR) is a promising topic in the recommender systems community, aiming to capture the users’ dynamic preferences and predict the next item they like based on their historical behaviors. Early studies mainly adopt Markov chains to model the transition patterns of users (Rendle et al., 2010). In recent years, due to the rapid development of deep learning, different kinds of neural networks have been introduced into the SR, which resulted in notable enhancements. The most representative works include GRU4Rec (Hidasi et al., 2015), Caser (Tang and Wang, 2018), SR-GNN (Wu et al., 2019), SASRec (Kang and McAuley, 2018), and BERT4Rec (Sun et al., 2019).
@@
With the thriving of PLMs, researchers begin to explore their application in SR. Some studies utilize PLMs as item text encoders, aiming to enhance item representations with the rich prior knowledge embedded within PLMs (Zhou et al., 2020; Mu et al., 2023; Wang et al., 2022b; Hou et al., 2022; Yuan et al., 2023). These methods always leverage a PLM to encode item text (e.g., title, brand) into text embeddings, which are then utilized to substitute or augment the original ID embeddings. Recently, propelled by the significant reasoning and long text modeling abilities demonstrated by Large Language Models (LLMs), some pioneering researchers have sought to harness the advanced sequence modeling capabilities of PLMs to enhance the performance of SR. One line of work uses fixed LLMs for recommendations via prompt or in-context learning, which achieve promising performance under the few-shot or zero-shot settings (Gao et al., 2023; Liu et al., 2023). However, due to the significant disparity between language and behavior modeling, these methods often perform poorly when interactions are sufficient. In that case, recent works choose to adopt recommendation objectives to fine-tune PLMs for adaptation, which achieve promising results (Geng et al., 2022; Cui et al., 2022a; Zhang et al., 2023; Li et al., 2023; Qu et al., 2023).
@@
Despite the efficacy of PLM-enhanced methods in existing literature, a critical question remains largely unexplored: do PLMs effectively and efficiently boost sequential recommendation by providing accurate prior knowledge and superior sequence modeling capabilities as theoretically expected? While some research has investigated their effect in item representation learning (Wang et al., 2022b; Yuan et al., 2023; Wei et al., 2023), their impact on behavior sequence modeling has yet to be thoroughly examined. Consequently, our study aims to figure out two questions: (1) whether the powerful sequence modeling capabilities of PLMs have been fully and economically utilized in the behavior sequence modeling of SR. (2) If not, is there a more effective, universal, and economical way for us to employ PLMs for SR?
@@
For question one, we conduct two motivating experiments (in Sec. 3) using a representative SOTA PLM-enhanced multi-domain SR model, RECFORMER (Li et al., 2023). The findings highlight two critical aspects: (1) RECFORMER functions differently from its base PLM backbone in behavior modeling. Its attention maps exhibit a clear functional stratification, with higher layers (more crucial for SR adaptation) more assimilated to conventional ID-based SR models. (2) There appears to be significant parameter redundancy when employing PLMs for behavior modeling. The global attentions of the sequence display considerable resemblance among different heads and layers in the same functional stratification. Only tuning 1/4 selected layers of RECFORMER can yield performance comparable to or even surpassing the original RECFORMER.
@@
Based on the above observation, we implement several enhanced variants of RECFORMER with simplified sequence modeling methods borrowed from classical ID-based SR models such as SASRec and BERT4Rec. We discover the key factor of improvement borrowed from PLM to SR: item representations built by behavior-tuned PLM. Experimental results indicate that PLM-involved SR models with simpler sequence models can achieve comparable results to RECFORMER with much better efficiency, suggesting that PLMs are not that suitable for behavior sequence modeling, for their superior sequence modeling and reasoning capacities in NLP are not fully leveraged and transferred into the task of SR. Consequently, we conclude a simple, effective, and universal framework to efficiently take advantage of PLM for SR: (a) adopting simplified (ID-based) sequential models for behavior sequence modeling (for question one), and (b) using behavior-tuned PLMs (rather than vanilla PLMs) for item embedding initialization (for question two). Further experiments demonstrate that our behavior-tuned PLMs initialized item embeddings are universal with different (ID-based or PLM-based) sequential models and settings. The contributions are concluded as follows:
@@
  • We introduce a novel framework that uses behavior-tuned PLMs for item initialization and simplified methods for behavior sequence modeling. Our framework is simple, effective, and universal, leading to significant improvements over PLM-based and classical ID-based SR baselines on different settings without additional inference costs.
@@ 2.1. Sequential Recommendation
Sequential recommendation (SR) aims to model the users’ dynamic behavior patterns based on their historical interactions. The very early works usually adopt Markov chains (MCs) to model the behavioral patterns of users. Rendle et al. (Rendle et al., 2010) integrates MF and MCs to simultaneously model users’ overarching preferences and their historical interactions. Beyond first-order MCs, several approaches have embraced higher-order MCs, incorporating more preceding items for sequential modeling (He et al., 2017; He and McAuley, 2016). Later, with the development of deep learning methods, various deep neural networks were introduced into SR. RNNs such as Gated Recurrent Unit (GRU) (Cho et al., 2014) are introduced to model sequential behaviors, including session-based GRU (e.g., GRU4Rec (Hidasi et al., 2015)) and user-based GRU (Donkers et al., 2017)). Moreover, CNNs have also been demonstrated to effectively capture short-term sequential dynamics using both horizontal and vertical convolutional filters (Tang and Wang, 2018). Many Graph Neural Network (GNN) based models (Wu et al., 2019; Guo et al., 2022; Wang et al., 2022a; Chang et al., 2021), were also implemented by researchers to seek better performance. Recently, due to the notable achievements in sequence modeling made by Transformers (Vaswani et al., 2017) in natural language processing, many attention-based models have also been intensively explored. SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) employ unidirectional and bidirectional Transformers separately for learning dynamic user preferences. LSSA (Xu et al., 2021) combines long-term and short-term self-attention mechanisms, designed to model users’ enduring preferences alongside their immediate needs. Self-supervised learning (SSL) is also introduced by recent works for improving the performance and training efficiency of SR models (Qiu et al., 2022; Chen et al., 2022; Xie et al., 2022; Wu et al., 2022; Zhou et al., 2020). Besides, some recent efforts explore the cross-domain sequential recommendation (Ma et al., 2019; Hao et al., 2021; Cao et al., 2022), while most of them face challenges in smoothly transferring the generalized information to new domains. To address this, some works (Wang et al., 2022b; Hou et al., 2022; Yuan et al., 2023) further use PLMs and vision encoders to fuse modality information into item representations for a more comprehensive understanding of item contents, functioning as the bridge across different datasets.
@@ 2.2. PLM for Recommendation
With the thriving of foundation models, many applications of PLMs for SR have been researched, hoping to leverage the rich knowledge and powerful sequence modeling capabilities of PLMs to enhance SR. Some works propose to use them as text encoders (e.g., build item representations from their textual representations based on PLMs rather than merely using the conventional random initialized ID embeddings). For example, IDA-SR (Mu et al., 2023) and UniSRec (Hou et al., 2022) apply BERT to obtain item representations from their corresponding texts. LLM2BERT4Rec (Harte et al., 2023) utilizes an embedding model with the titles of the items to retrieve their embeddings. Besides, some researchers try to convert users’ historical interactions and items into plain text and utilize them to fine-tune the PLMs for SR (Qu et al., 2023; Cui et al., 2022b; Li et al., 2023; Geng et al., 2022).
Recently, the emergence of Large Language Models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Ouyang et al., 2022) has revolutionized the field of natural language processing with their powerful reasoning abilities and rich world knowledge. Therefore, some pioneer works explore the possibility of directly using powerful LLMs for recommendation via prompt or in-context learning (Hou et al., 2023; Gao et al., 2023; Wang and Lim, 2023; Liu et al., 2023; Kang et al., 2023). Although these methods achieve impressive results in cold-start scenarios, they still perform poorly in SR when interactions are sufficient since behavioral information is essential for making precise recommendations. Another line of work considers introducing collaborative signals into LLMs to improve their recommendation performance. TallRec (Bao et al., 2023b) and InstructRec (Zhang et al., 2023) perform instruct-tuning on LLMs to adapt them for understanding behavioral patterns. RLMRec (Ren et al., 2023) aligns the semantic space of LLMs with collaborative relational signals through cross-view alignment to improve the quality of representations. BIGRec (Bao et al., 2023a) treats SR as a text generation task, generating item titles by a two-step approach with the LLM fine-tuned on user behavioral sequences. Nevertheless, these methods lack a thorough examination of whether PLMs’ capabilities are being effectively and efficiently harnessed. In our paper, we deeply explore the involvement of PLMs in SR tasks. We intend to scrutinize whether their formidable abilities are truly being utilized to their full potential and offer insights on how to make the most out of PLMs in the context of SR.
@@ 3. Analyses on Behavior Sequence Modeling of PLM-based SR Models
In this section, we comprehensively analyze the capability and suitability of PLM-based SR models on behavior sequence modeling. Without loss of generality, we employ the SOTA PLM-based SR model RECFORMER (Li et al., 2023) for the following explorations.
@@ 3.1. Background of RECFORMER
RECFORMER is a PLM-based SR model that comprises a classical PLM Longformer (Beltagy et al., 2020) and two additional embeddings that indicate the type of token (determining the token belongs to the special token [CLS], textual attribute key or textual attribute value) and its associated item (which item the token belongs to). It transforms items and behavioral sequences into text sequences and then utilizes the behavior-tuned model to obtain their corresponding representations for providing recommendations. RECFORMER’s training process has three steps, including pre-training and two-stage fine-tuning. To simplify, we designate these stages as stage-PT, stage-FT1, and stage-FT2 respectively. (1) In stage-PT, RECFORMER adopts Masked Language Modeling (MLM) and item-item contrastive (IIC) tasks to pre-train the model. Considering the computational costs, RECFORMER treats in-batch next items as negative instances instead of negative sampling or fully softmax for the IIC task. (2) In stage-FT1, RECFORMER encodes all items before each epoch starts, obtaining their corresponding item embeddings, then fine-tuning all parameters with these frozen embeddings (for efficiency considering the large size of item candidates). (3) In stage-FT2, the RECFORMER selects the optimal model from stage-FT1 to acquire its item embeddings and fix them. Stage-FT2 only tunes the parameters of the behavior sequence modeling part (similar to stage-FT1, but with the item embeddings remaining fixed and not updated each epoch). This strategy of maintaining fixed item embeddings is primarily motivated by the significant computational demands associated with generating all item representations during the cross-entropy loss calculation. Therefore, the fine-tuning process of RECFORMER could be simplified as: (1) obtaining the item embeddings from behavior-tuned Longformer, and (2) tuning PLM solely for behavior sequence modeling with fixed item embeddings. The latter step is our focus.
@@ 3.2. Analysis on Attention Discrepancy
Intuitively, the sequential patterns of items in SR and tokens in LM significantly differ. SR tends to prioritize the most recent interactions of users, focusing more on their immediate past behavior to make predictions (Kang and McAuley, 2018; Sun et al., 2019). On the other hand, LM concentrates more on maintaining linguistic consistency and adhering to the syntactic structure of the text, ensuring coherence and grammatical accuracy throughout the content. To verify this assumption, we visualize the attention distributions of the special token [CLS] that is used to get the user representation for recommendation in RECFORMER (and its backbone Longformer (Beltagy et al., 2020) for comparisons).
Specifically, we follow the setting in (Li et al., 2023) to pre-train and fine-tune RECFORMER, and stack the attention scores from different heads in the same layer on the vertical axis. As for Longformer, we directly utilize its original checkpoint, drawing the attention maps in the same way as RECFORMER. It’s important to highlight that RECFORMER reverses the user’s interaction sequence. Consequently, within the attention maps, items positioned to the left symbolize those with which the user has most recently interacted. In Fig. 1, we provide the attention maps of two instances from Pantry and Instruments (Ni et al., 2019) with its length of historical sequence as three (for better visualization). Upon examining various instances from multiple domains, we have some universal findings: